### v2

@@ 4 Network Architecture
In this section, we describe the network architecture used in our experiments. In order to use an architecture that matches our theoretical findings exactly we use a DeepONet architecture as seen in Figure 2. The input for the Branch network are the measurements of the EIT experiment measurements. That is an array of $208$ measurements. This number appears because we using $16$ different current setups of adjacent electrodes. For each current setup, consisting of one adjacent source and sink electrode we measure the voltage differences (adjacent electrodes) on the other $14$ electrodes. Hence, we obtain $13$ measurements per current setup. In total, we get $16*13=208$ measurements. We then use a series of fully connected ResNet blocks to output the same dimensionality as the later defined Trunk network (See Table LABEL:BranchNet).
@@
One interesting observation from the computational perspective is that, for each EIT experiment, the Branch network only has to be computed once. The Trunk network has to be evaluated for many points (in parallel). That is why the Trunk network is chosen to be much more lightweight than the Branch network. We note that we experimented and evaluated our method with many different network architectures and that from the set of DeepONet, this setup seemed to perform particularly well.
@@ 5.1.1 Training Data
During training we generate a point cloud 1024 points $(x,y)\in\mathbb{R}^{2}\subset\Omega$ and the corresponding $\sigma_{x,y}$. Our sampling is to be understood as sampling $50\%$ of the points from the background and $50\%$ of the points from the anomalous regions. This is only done if there are anomalies in $\sigma$. The reason for this kind of sampling is that in our experience it significantly increased the correct detection on small anomalies. The idea behind this is that if only a very small amount of samples is anomalous the neural network has the tendency of predicting everything as background.
@@ 5.1.2 Training and Testing
Our neural network is of the form shown in Figure 2, consisting of the two components defined in Table LABEL:BranchNet and LABEL:TrunkNet. During the Training process, we first evaluate the measurement data via our Branch network followed by evaluating the corresponding large set of 1024 $2D$ points with the Trunk network. Then the inner product of both outputs is computed. Each estimated value $\hat{\sigma}_{x,y}$ is then compared with the ground truth $\sigma_{x,y}$ with an $L^{2}$-loss. Note that because we use $50\%$ samples of the anomalous area and the rest of the background area this translates to a weighted $L^{2}$-loss where the anomalous area has significantly more weight than it would naturally have if we were comparing images on a PDE mesh. During the Training process, we use a Batch size of 32 conductivity setups. The training is performed for 200 epochs using the ADAM optimizer.
We note that in the 1-D Cut plot in Figure 3 we post processed $\sigma_{x,y}$ at point $(x,y)$ if $\sigma_{x,y}$ was below a threshold. In those cases we set $\sigma_{x,y}$ to the threshold. We used the same method and threshold for both the IRGN baseline and the DeepONet.
@@ 5.2 Performance
We compare our method with IRGN, a numeric state-of-the-art gradient-based optimization method. We do not provide an extensive comparison with other methods because our main results are on the theoretical side. Hence, the goal is to show that our method is competitive with one state-of-the-art method.
5.2.1 Baseline: Iteratively Regularized Gauss-Newton Method (IRGN)
We use the Iteratively Regularized Gauss-Newton (IRGN) method as described in [3] as a baseline for our approach. The IRGN method is a state of the art method for solving the EIT problem in an analytical setting on a PDE mesh. We use the standard Tikhonov regularized IRGN method, see [3] for a good description of this method.
@@ 5.2.2 Comparison
We evaluate both methods on 5 images at $1\%$ additive measurement noise. The comparison results can be found in Figure 3. On the left-hand side is the ground truth image. The second row is the IRGN method followed by our proposed method. Note the first three figures are in the same color range. The rightmost image is a $1D$ cut through the previous 3 images. The cut direction is visualized in a white line in the ground truth image. We observe that both methods are approximately finding the correct location of the anomaly. However, our method outperforms the IRGN method in terms of obtaining the correct sigma in the anomalous areas. Some of the numerical values can be found in noise study in Table LABEL:noise_table
@@ 5.3 Noise Study
The stability of the method against higher noise levels on a single image is evaluated. All experiments of the errors are repeated ten times (On Fig. 3, Row 1) and the mean and standard deviation are displayed in Table LABEL:noise_table. Overall, the $L^{2}$-Error is increasing in a noise pattern in the DeepONet method. The same is true for the IRGN method. However, the IRGN method does have lower $L^{2}$-errors.
We believe that the $L^{2}$ error is not the correct metric because, for images with small anomalous regions, a small error does not imply that the location and conductivity value of the anomaly were estimated properly. For example, in such a case, the reconstruction where everything is exactly the background value, would have relatively low $L^{2}$ errors. However, the reconstruction would not have any practical use. Hence, we use a weighted $L^{2}$ error that takes $50\%$ weight of the ground truth anomalous areas and the rest from the background area. In this metric, the DeepONet method outperforms the IRGN method as expected from seeing the images.
The same patterns in both the weighted and unweighted $L^{2}$ errors are found in all other images that we tested. We note that for large noise levels from $7\%$, the reconstructed anomalies find significantly worse locations and shapes.
Table 3: Noise Study - Fig. 3, Row 1
Noise Level $1\%$ $2\%$ $3\%$ $4\%$ $5\%$ $9\%$
DeepONet - $L^{2}$ Error $1.53\pm 0.14$ $1.75\pm 0.31$ $1.78\pm 0.34$ $2.27\pm 0.44$ $2.46\pm 0.85$ $3.86\pm 1.17$
IRGN - $L^{2}$ Error $0.37\pm 0.02$ $0.42\pm 0.03$ $0.53\pm 0.07$ $0.64\pm 0.11$ $0.81\pm 0.20$ $1.64\pm 0.36$
DeepONet - Weighted $L^{2}$ $0.54\pm 0.52$ $0.63\pm 0.66$ $0.65\pm 0.70$ $0.86\pm 0.96$ $1.05\pm 1.25$ $1.62\pm 2.17$
IRGN - Weighted $L^{2}$ $14.16\pm 0.79$ $14.15\pm 0.98$ $14.34\pm 2.87$ $13.44\pm 2.61$ $14.65\pm 3.75$ $14.21\pm 4.71$
@@ 6 Conclusion
In this work, we have generalized the earlier known approximation-theoretic results for approximating a map between infinite-dimensional function spaces using DeepONets to the case where we can provide similar guarantees even when the goal is to learn a map between a space of (N-t-D) operators to the space of conductivities (functions). This map arises naturally in the framework of Electrical Impedance Tomography which is a non-invasive medical imaging modality that is becoming increasingly prominent in especially several neuro-imaging applications. As deep-learning networks are becoming increasingly popular, it behooves us to go beyond black-box description of such networks and to make them interpretable so as to facilitate deployment with confidence in critical applications such as medical imaging. Providing approximation-theoretic guarantees of the kind given in this work is a crucial step in achieving the goal of interpretable neural-networks. In fact, we believe that the analysis carried out in this work can be easily adapted to many other problems wherein the input and output spaces are classes of operators satisfying certain general assumptions. To that end, we have proposed a DeepONet architecture that is motivated by the underlying inverse map appearing in EIT. We have compared our approach against the widely used IRGN benchmark and show that the proposed network outperforms the baseline in numerical studies. We also performed a noise study and show that robustness of the deep-learning based against a weighted $L^{2}$ error metric as introduced in this work. To our knowledge, this is the first work that uses a DeepONet architecture for approximating an operator-to-function map as appearing in EIT, while also providing a theoretical guarantee to show its effectiveness. As already stated, our numerical results back-up our theoretical findings.
@@ 7 Appendix
Lemma 3.
@@
Proof.
@@
First of all we note that $\mathbb{H}_{r}$ is a separable Hilbert Space with a countable basis given in [1, section 5.1]. From [63, Theorem 197], it suffices to show that $D_{\Lambda}$ is closed, bounded and has equi-small tails with respect to the orthonormal basis for $\mathbb{H}_{r}$. The fact that $D_{\Lambda}$ is bounded follows from [20, (2.5)] as before. The fact that $D_{\Lambda}$ has equi-small tails follows from [1, Lemma 18] in which we choose $J,K$ and $\nu$ such that $C\min(J,K)^{-\nu}\leq\epsilon$ for any given $\epsilon$. Thus it remains to show that $D_{\Lambda}$ is closed. To that end, choose a sequence of relative N-t-D operators $(\tilde{\Lambda}_{\gamma_{n}}:\gamma_{n}\in\Gamma)$ such that they converge to some operator $T$ in $\mathbb{H}_{r}$ i.e., $\tilde{\Lambda}_{\gamma_{n}}\to T$. We will show that there exists some $\gamma\in\Gamma$ such that $T=\tilde{\Lambda}_{\gamma}$ i.e. $T$ is also a N-t-D operator. First of all note that since $(\tilde{\Lambda}_{\gamma_{n}})$ converges, hence $(\tilde{\Lambda}_{\gamma_{n}})$ is actually Cauchy in the $\mathbb{H}_{r}$ norm. Mimicking the proof of [1, Lemma 5], we can show that $(\Lambda_{\gamma_{n}})$ is also Cauchy in the operator norm $\mathcal{L}(L^{2}\to L^{2})$. By using the inverse continuity result for the Calderon problem, see [27, Theorem 2.3], we get that the sequence $\gamma_{n}$ is Cauchy with respect to the $L^{\infty}$ norm and hence converges, say $\gamma_{n}\to\gamma_{0}\in L^{\infty}(\Omega)\cap\Gamma$. Let us now consider, $\Lambda_{\gamma_{0}}$. From forward continuity result for the Calderon problem, see [27, Lemma 2.5], $\Lambda_{\gamma_{n}}\to\Lambda_{\gamma_{0}}$. Now from uniqueness of the limit, $\Lambda_{\gamma_{0}}=T$. Hence the set, $D_{\Lambda}$ satisfies all theree criteria for it be a compact subset of $\mathbb{H}_{r}$. ∎
Lemma 4.
@@
[16, Theorem 4.1] Let $X$ be an arbitrary metric space, $A$ a closed subset of $X$, $L$ a locally convex linear space, and $f:A\to L$ be a continuous map. Then there exists an extension $F:X\to L$ (which is continuous by definition) such that $F(a)=f(a)$ for all $a\in A$.
7.1 Continuity results of some maps
@@
Consider the encoding map $\mathcal{E}(\tilde{\Lambda}):=\mathcal{E}_{M}(\tilde{\Lambda})=\tilde{R}^{M}_{\gamma}.$ We have the following continuity results:
Lemma 5.
@@
[44, section 2] The map $T:\gamma\to\tilde{R}_{\gamma}^{M}$ is Lipschitz continuous, i.e. $\lvert\lvert\tilde{R}^{M}_{\gamma_{1}}-\tilde{R}^{M}_{\gamma_{2}}\rvert\rvert_{\mathcal{L}(L^{2}\to L^{2})}\leq\lvert\lvert\gamma_{1}-\gamma_{2}\rvert\rvert_{L^{\infty}}$
Lemma 6.
@@
[27, Theorem 2.3] $\lvert\lvert\gamma_{1}-\gamma_{2}\rvert\rvert_{L^{2}}\leq c\lvert\lvert\gamma_{1}-\gamma_{2}\rvert\rvert_{L^{\infty}}\leq C\lvert\lvert\tilde{\Lambda}_{\gamma_{1}}-\tilde{\Lambda}_{\gamma_{2}}\rvert\rvert_{\mathcal{L}(L^{2}\to L^{2})}$

### v3

@@ 4 Network Architecture
In this section, we describe the network architecture used in our experiments which corresponds to the one depicted in Fig. 2. The input for the Branch network are the measurements of the EIT experiment. For the numerical simulations, we imagine using 16 different electrodes through which the current will be injected and the voltages measured. In practical settings, current is injected by using a pair of adjacent source and sink electrodes. For each such current pattern, voltage differences are measured across adjacent electrodes on the remaining $14$ electrodes. Hence, we obtain $13$ voltage difference measurements per injected current pattern. In total, we get $16*13=208$ measurements which can be arranged in the form of a matrix. Many a times, one uses zero-padding to fill in the missing entries of the matrix to make it a square matrix. The measurements are then fed into a series of fully connected ResNet blocks to output a vector of the same dimensionality as the output of the later defined Trunk network (See Table LABEL:BranchNet).
@@
One interesting observation from the computational perspective is that, for each EIT experiment, the Branch network only has to be computed once. The Trunk network has to be evaluated for many points (in parallel). That is why the Trunk network is chosen to be much more lightweight than the Branch network. We note that we experimented and evaluated our method with many different network architectures of similar sizes and they all performed reasonably well. We chose to present this setup as this seemed to do the best in our experiments.
@@ 5.1.1 Training Data
During training we generate a point cloud 1024 points $(x,y)\in\mathbb{R}^{2}\subset\Omega$ and the corresponding $\sigma_{x,y}$. Our sampling is to be understood as sampling $50\%$ of the points from the background and $50\%$ of the points from the anomalous regions. This is only done if there are anomalies in $\sigma$. The reason for this kind of sampling is that in our experience it significantly increased the correct detection on small anomalies. The idea behind this is that if only a very small amount of samples is anomalous the neural network has the tendency of predicting everything as background. We note that more sophisticated sampling methods have been advocated in the literature. However, for the verification of our theoretical results, this simple sampling scheme tends to also work well.
@@ 5.1.2 Training and Testing
Our neural network is of the form shown in Figure 2, consisting of the two components defined in Table LABEL:BranchNet and LABEL:TrunkNet. During the Training process, we first evaluate the measurement data via our Branch network followed by evaluating the corresponding large set of 1024 $2D$ points with the Trunk network. Then the inner product of both outputs is computed. Each estimated value $\hat{\sigma}_{x,y}$ is then compared with the ground truth $\sigma_{x,y}$ with a weighted- $L^{2}$-loss. The weighted $L^{2}$ loss is used to reflect the fact that during training we trained our network using $50\%$ points from the anomalous area and the rest from the background area. Furthermore, during the training process, we use a Batch size of 32 conductivity setups. The training is performed for 200 epochs using the ADAM optimizer. The 1-D cut plots in Figure 3 denotes the values of conductivity function along a 1-D profile indicated by a white line in the original phantom.
@@ 5.2 Performance
In this subsection, we describe the performance of our proposed method in our numerical studies and compare it against a standard method used for EIT inversion.
5.2.1 Baseline: Iteratively Regularized Gauss-Newton Method
We use the Iteratively Regularized Gauss-Newton (IRGN) method, see e.g. [3], as a baseline for our approach. The IRGN method is a popular gradient-based optimization method for solving the EIT problem in an analytical setting on a PDE mesh. We use the standard Tikhonov regularized IRGN method, see [3] for a good description of this method. We do not provide an extensive comparison with other methods because our main results are on the theoretical side. Hence, the goal is to show that our method is competitive with one state-of-the-art method.
@@ 5.2.2 Comparison
We evaluate both methods on 5 images at $1\%$ additive measurement noise. The comparison results can be found in Figure 3. On the left-hand side is the ground truth image. The second row is the IRGN method followed by our proposed method. Note the first three figures are in the same color range. The rightmost image is a $1D$ cut through the previous 3 images. The cut direction is visualized in a white line in the ground truth image. We observe that both methods are approximately finding the correct location of the anomaly. However, our method outperforms the IRGN method in terms of obtaining the correct sigma in the anomalous areas. Some of the numerical values can be found in noise study in Table LABEL:noise_table.We also note the good generalization ability of the proposed network in practice. In particular, observe that in the third and fourth row of Fig 3, we attempt to reconstruct thin and stretched rectangular shaped anomalies, while the network was trained only on circular anomalies.
@@
5.3 Weighted $L^{2}$ error
We believe that the standard $L^{2}$ error is not the correct metric because, for images with small anomalous regions, a small error does not imply that the location and conductivity value of the anomaly were estimated properly. For example, in such a case, the reconstruction where everything is exactly the background value, would have relatively low $L^{2}$ errors. However, the reconstruction would not have any practical use. Hence, we use the weighted $L^{2}$ error that takes $50\%$ weight of the ground truth anomalous areas and the rest from the background area. In this metric, the DeepONet method outperforms the IRGN method as expected from seeing the images as well as the error table below.
@@ 5.4 Noise Study
The stability of the method against higher noise levels on a single image is evaluated. All experiments of the errors are repeated ten times (On Fig. 3, Row 1) and the mean and standard deviation are displayed in Table LABEL:noise_table. Overall, the weighted $L^{2}$-Error is increasing in a noise pattern in the DeepONet method. The same is true for the IRGN method. However, the IRGN method does have lower errors. We note that for large noise levels from $7\%$, the reconstructed anomalies find significantly worse locations and shapes.
Table 3: Noise Study - Weighted $L^{2}$ Fig. 3, Row 1
Noise Level $1\%$ $2\%$ $3\%$ $5\%$ $9\%$
DeepONet $0.54\pm 0.52$ $0.63\pm 0.66$ $0.65\pm 0.70$ $1.05\pm 1.25$ $1.62\pm 2.17$
IRGN $14.16\pm 0.79$ $14.15\pm 0.98$ $14.34\pm 2.87$ $14.65\pm 3.75$ $14.21\pm 4.71$
@@ 6 Conclusion
In this work, we have generalized the earlier known approximation-theoretic results for approximating a map between infinite-dimensional function spaces using DeepONets to the case where we can provide similar guarantees even when the goal is to learn a map between a space of (N-t-D) operators to the space of conductivities (functions). This map arises naturally in the framework of Electrical Impedance Tomography which is a non-invasive medical imaging modality that is becoming increasingly prominent in especially several neuro-imaging applications. As deep-learning networks are becoming increasingly popular, it behooves us to go beyond black-box description of such networks and to make them interpretable so as to facilitate deployment with confidence in critical applications such as medical imaging. Providing approximation-theoretic guarantees of the kind given in this work is a crucial first step in achieving the goal of interpretable neural-networks. In fact, we believe that the analysis carried out in this work can be easily adapted to many other problems wherein the input and output spaces are classes of operators satisfying certain general assumptions. We have compared our approach against the widely used IRGN benchmark and show that the proposed network outperforms the baseline in numerical studies. We also performed a noise study and show that robustness of the deep-learning based against a weighted $L^{2}$ error metric as introduced in this work. To our knowledge, this is the first work that uses a DeepONet architecture for approximating an operator-to-function map as appearing in EIT, while also providing a theoretical guarantee to show its effectiveness. As already stated, our numerical results back-up our theoretical findings.
@@ 7 Appendix
In this section, we collect some auxiliary results needed for the proof of our main theorem.
Lemma 7.1.
@@
Proof 7.2.
First of all we note that $\mathbb{H}_{r}$ is a separable Hilbert Space with a countable orthonormal basis given in [1, section 5.1]. We recall the following theorem which is given in e.g. [15, Chapter 2], [65, Theorem 197].
Theorem 7.3.
[65, Theorem 197] Let $H$ be a separable Hilbert space, and let $\{e_{k}\}_{k=1}^{\infty}$ be an orthonormal basis of $H$. Then a subset $K\subset H$ is compact iff:
  1. 1.
    $K$ is closed,
  2. 2.
    $K$ is bounded, and
  3. 3.
    $K$ has equi-small tails with respect to $\{e_{k}\}$, i.e., for every $\varepsilon>0$, there exists $N\in\mathbb{N}$ such that for all $\vec{v}\in K$,
    $\sum_{k=N}^{\infty}|\langle\vec{v},e_{k}\rangle|^{2}<\varepsilon^{2}.$
@@
Now, the fact that $D_{\Lambda}$ is bounded follows from [21, (2.5)] as before. The fact that $D_{\Lambda}$ has equi-small tails follows from [1, Lemma 4] in which we choose $J,K$ and $\nu$ such that $C\min(J,K)^{-\nu}\leq\varepsilon$ for any given $\varepsilon$. Thus it remains to show that $D_{\Lambda}$ is closed. To that end, choose a sequence of relative N-t-D operators $(\tilde{\Lambda}_{\gamma_{n}}:\gamma_{n}\in\Gamma)$ such that they converge to some operator $T$ in $\mathbb{H}_{r}$ i.e., $\tilde{\Lambda}_{\gamma_{n}}\to T$. We will show that there exists some $\gamma\in\Gamma$ such that $T=\tilde{\Lambda}_{\gamma}$ i.e. $T$ is also a N-t-D operator. First of all note that since $(\tilde{\Lambda}_{\gamma_{n}})$ converges, hence $(\tilde{\Lambda}_{\gamma_{n}})$ is actually Cauchy in the $\mathbb{H}_{r}$ norm. Using the relation between the operator norm and the information-theoretic Hilbert-Schmidt norm as stated in [1, Lemma 5], we can show that $(\Lambda_{\gamma_{n}})$ is also Cauchy in the operator norm $\mathcal{L}(L^{2}\to L^{2})$. By using the inverse continuity result for the Calderon problem, see [28, Theorem 2.3], we get that the sequence $\gamma_{n}$ is Cauchy with respect to the $L^{\infty}$ norm and hence converges, say $\gamma_{n}\to\gamma_{0}\in L^{\infty}(\Omega)\cap\Gamma$. Let us now consider, $\Lambda_{\gamma_{0}}$. From forward continuity result for the Calderon problem, see [28, Lemma 2.5], $\Lambda_{\gamma_{n}}\to\Lambda_{\gamma_{0}}$. Now from uniqueness of the limit, $\Lambda_{\gamma_{0}}=T$. Hence the set, $D_{\Lambda}$ satisfies all three criteria for it be a compact subset of $\mathbb{H}_{r}$.
Lemma 7.4.
@@
[17, Theorem 4.1] Let $X$ be an arbitrary metric space, $A$ a closed subset of $X$, $L$ a locally convex linear space, and $f:A\to L$ be a continuous map. Then there exists an extension $F:X\to L$ (which is continuous by definition) such that $F(a)=f(a)$ for all $a\in A$.
@@
Finally, consider the encoding map $\mathcal{E}(\tilde{\Lambda}):=\mathcal{E}_{M}(\tilde{\Lambda})=\tilde{R}^{M}_{\gamma}.$ We have the following continuity results:
Lemma 7.5.
@@
[45, section 2] The map $T:\gamma\to\tilde{R}_{\gamma}^{M}$ is Lipschitz continuous, i.e. $\lvert\lvert\tilde{R}^{M}_{\gamma_{1}}-\tilde{R}^{M}_{\gamma_{2}}\rvert\rvert_{\mathcal{L}(L^{2}\to L^{2})}\leq\lvert\lvert\gamma_{1}-\gamma_{2}\rvert\rvert_{L^{\infty}}$
Lemma 7.6.
@@
[28, Theorem 2.3] $\lvert\lvert\gamma_{1}-\gamma_{2}\rvert\rvert_{L^{2}}\leq c\lvert\lvert\gamma_{1}-\gamma_{2}\rvert\rvert_{L^{\infty}}\leq C\lvert\lvert\tilde{\Lambda}_{\gamma_{1}}-\tilde{\Lambda}_{\gamma_{2}}\rvert\rvert_{\mathcal{L}(L^{2}\to L^{2})}$